OpenStudy (anonymous):

I am lost on this how do you figure out this equation in inches: (6 7/8)x (7 3/4)(21 3/5) ------- = ------------- 6 7/8 6 7/8

OpenStudy (anonymous):

you might want try retyping this....

OpenStudy (anonymous):

(6 7/8)x = (7 3/4)(21 3/5) ------- ------------- 6 7/8 6 7/8

OpenStudy (anonymous):

Here is the answer x~ 24 2/5 inches (The answer was to be simplify. Round to the nearest 1/5 inch) I have know clue how they come up 24 2/5 inches

OpenStudy (anonymous):

since the denominator are the same, they cancel, so you're left with (6 7/8)x = (7 3/4)(21 3/5) change everything to improper (55/8)x = (31/4)(108/5) check my calculation (55/8)x = 837/5 x = (837/5)*(8/55) x = 6696/275 ....definitely check my calculations..

OpenStudy (anonymous):

oh wait you gave me the answer yeah x = 6696/275 ~ 24.3490

OpenStudy (anonymous):

Where does the 837/5 come from?

OpenStudy (anonymous):

\[\frac{31}{4} * \frac{108}{5} = \frac{837}{5}\] look at the link above it

OpenStudy (anonymous):

Then how do you figure this out: Suppose to round to the whole centimeter (6 7/8)y = ( 7 3/4)(58) ------- --------- 6 7/8 6 7/8

OpenStudy (anonymous):

The answer is 62 centimeters

OpenStudy (anonymous):

I figure that one out but how would you do this one (7 1/4) y = (7 1/8)58) ------- --------- 7 1/4 7 1/4

OpenStudy (anonymous):

(7 1/4) y = (7 1/8)(58) (7 1/4) y = (413 1/4) y = 57

OpenStudy (anonymous):

I hate to be stupid but how did you get 57?

OpenStudy (anonymous):

(7 1/4) y = (7 1/8)(58) ^----^--- multiply together, you will get 413 1/4 (7 1/4) y = (413 1/4) y = (413 1/4) / (7 1/4) y = 413 / 7 y = 59 <-- correction, i meant to type 59 i'm sorry